A car accelerates from rest at constant rate of 2 m/sec/sec for some time. Then it retards at a constant at a constant rate of 4 m/sec/sec and comes to rest. What is the maximum speed attained by the car, if it remains in motion for 3 seconds?


options: (1) 2 m/sec  (2) 3 m/sec (3) 4 m/sec (4) 6 m/sec


pls give explanation also...u may reply on my email ID... anurag_can@yahoo.com....also which is best textbook for Kinematics? thanks

Grade:12

1 Answers

Prateek Gupta
11 years ago

I think the ans is 4m/sec...Draw V-t Graph properly and then u can solve

use the fact the slope of vt graph is acceleration and slope = tan theta = Perpendicular(here maximum velocity)/Base

Let Times be t1 and t2

acc. to graph ull get 2=P/t1  and 4=P/t2

or we can write 1/2 = t1/P   and 1/4 = t2/P

adding both we will get t1 + t2/P or 3/p = 3/4                      (t1 + t2 = 3)

SOLVE AND FIND p..

..I am too an iit aspirant +1 class..i THINK cS PANDEY HC VERMA ARE GOOD BUS TO START WITH..;)